The property of oxygen to dissolve into water is very important to aquatic environmental systems The following Henrys constants are found in the literature for fresh water: Temperature (C KH (mol-L atm1) 10 15 20 25 1.90x103 1.68x10-3 1.50x103 1.35x103 1.22x103 Develop a spreadsheet to calculate and plot the saturation concentration (in mg/L) of oxygen (02) in fresh water. Assume the atmospheric pressure is 1 atm and the mole fraction of oxygen in the air is 0.21.
Answer #1

Henry's law is given by

Yi*P= Hi*Ci ....................1

where Yi is mole fraction in vapor phase

P = total pressure of the system

Hi = Henry's constant

Ci is the mole fraction in liquid phase

Yi = 0.21

P = 1 atm

substituting all the values in equation (1)

0.21*1 = Xi*C

C = 0.21/H

now we have H for different temperature, now we will find C for different values of H and represent in the following table

T (oC) Kh(mol L-1 atm-1) Xi= 0.21/H C (mg/L)
5 0.0019 110.5263158 110.5263
10 0.00168 125 125
15 0.0015 140 140
20 0.00135 155.5555556 155.5556
25 0.00122 172.1311475 172.1311